# The Less Intelligent the Elements, the More Intelligent the Whole. Or, Possibly Not?

### v4

@@ 3.1 It’s Difficult to Leave a KISS
As illustrated above, the basic Wilensky and Reisman’s model does not support $E_{3}$. Suppose that predators and preys devise a reproductive strategy which, in their intentions, should be able to keep the system at $E_{3}$. With a slight departure from pure KISS, they may imagine that a flexible and adaptive strategy based on some negative feedback may generate co-existence. For instance, in order not to be either too many when the other species shrinks, or too few when the other species thrives, either predators, or preys, or both of them may decide to bind their reproduction to the other species’s success:
@@
Thus, this experiment apparently confirms Proposition 1, namely that individual intelligence destroys collective intelligence. However, these results were obtained with individuals who were not even attempting to figure out the global consequences of their behaviour. It was a very slight departure from pure KISS. Perhaps, meta-cognition — the maximum KIDS of Proposition 3 — can help devising individual behavior that generates co-existence.
3.2 The Extreme KIDS of Good Guys
Let us now explore the possibility that, as expressed by Proposition 3, at least some individuals can envision collective dynamics and therefore are able to influence it by selecting some appropriate behavior. Specifically, let us suppose that either predators, or preys, or both of them take on a collaborative attitude that may favor a collective agreement as prescribed by Qualification 3.3.
@@ 3.2 The Extreme KIDS of Good Guys
Finally, these behaviors can be combined by assuming that both predators and preys stop reproducing if their own population becomes larger than the other one. While the previous decisions were the outcome of agreements between either predators or preys, the combined behavior could be the outcome of a general agreement involving both populations.
@@
If predators behave as in 4, $E_{1}$ is reached only $2.8\%$ of the times, $E_{2}$ just $0.5\%$ of the times, whereas $E_{3}$ is reached $96.7\%$ of the times. Thus, this behavior succeeds to make predators and preys co-exist. To a closer scrutiny, this $96.7\%$ arises out of a $43.7\%$ of outcomes where the two populations are roughly constant, a $13.2\%$ where the population of preys oscillates whereas predators do not, a $21.0\%$ where the population of predators oscillates whereas preys do not and a $18.8\%$ where both populations oscillate.
@@
If both predators and preys take on the collaborative attitude 6 the outcome is somewhat mixed, with $E_{1}$ nonexistent at $0.0\%$ whereas both $E_{2}$ and $E_{3}$ are quite substantial at $22.9\%$ and $77.1\%$, respectively. Co-existence is the most likely outcome but happens less often than in 4. Unlimited growth of preys is substantial but far from inevitable as in case 5.
Here it is worth noticing that, in the Lotka-Volterra model, predators and preys do not enjoy the same status. On the one hand, predators capable of meta-cognition are in position to steer the ecology towards co-existence by exerting double-loop learning (Proposition 3.1). On the other hand, preys capable of meta-cognition understand the system but cannot steer it, as it is typical of deutero-learning (Proposition 3.2). Preys can at most obtain dominance in the short run, but at the cost of unsustainable explosive dynamics for themselves.
@@
3.3 The Extreme KIDS of Realpolitik
The problem with behavioral rules 4, 5 and 6 is that either preys or predators, or both, are supposed to be capable of sacrificing their immediate interests. In practice, this may be difficult to attain.
Is it possible to obtain stable coexistence of species without such a high degree of public spirit? We still want individuals to use their intelligence to understand the system’s global dynamics, but we also want them to pursue their own interests.
Proposition 2 can be combined with Proposition 3 to endow our agents with the ability to make predictions based on global understanding of the Lotka-Volterra model that they inhabit. Specifically, predators predict their own extinction if they either see the population of preys or their own population growing very quickly, whereas preys predict their own extinction if they see predators growing very quickly.
If they are rational, neither predators nor preys wait for their Lotka-Volterra world to unfold their inevitable destiny. As Qualification 2 suggests, they collectively agree to commit suicide if they see their inevitable end forthcoming. Thus, let us explore the following behavioural rules:
@@ 3.3 The Extreme KIDS of Realpolitik
These behavioural rules require an additional parameter to specify what “much faster” means. We introduced this parameter with the default assumption that “much faster” means ten times faster, supporting this assumption with a sensitivity analysis reported in Appendix C.
@@
Notably, by assuming that either predators or preys, or both of them are capable to predict the correct and inevitable outcome of the Lotka-Volterra model, co-existence $E_{3}$ can never be reached. Apparently, one cannot escape KISS (Proposition 1).
@@ 3.4 Prediction based on Extrapolation
Similarly to§ 3.3, also in this case it is necessary to add a parameter in order to specify how many past steps the above variations are computed on. We used this parameter with the default value of five steps, supporting this choice with a sensitivity analysis reported in Appendix C.
Behaviour 10 is the second most effective in generating co-existence after 4 (altruistic predators seeking a collective agreement, § 3.2). In particular, with 10 generalized extinction $E_{1}$ never occurs. Preys-dominated equilibrium $E_{2}$ is reached $8.4\%$ of the times and, most importantly, $E_{3}$ is reached $91.6\%$ of the times. Out of this $91.6\%$, in $7.7\%$ both populations oscillate, $78.1\%$ only predators oscillate, in just $0.4\%$ of cases the opposite happens and, finally, in the remaining $5.4\%$ neither population oscillates. In a nutshell, the most likely outcome is that the population of predators oscillates while the population of preys remains roughly constant.
@@
Interestingly, the mixed arrangement 12 is not some sort of average between 10 and 11 but rather a case on its own. With 12, $E_{1}$ is reached $17.9\%$ of the times, $E_{2}$ a mere $2.4\%$ whereas $E_{3}$ is reached $79.7\%$ of the times. Out of this $79,7\%$, $61.7\%$ of the times both populations oscillate, $1.4\%$ only the population of predators oscillates, $14.2\%$ the opposite happens and, finally, in the remaining $2.4\%$ neither population oscillates. Thus, co-existence is reached quite often, and most of the times it is reached with both populations oscillating.
This sort of co-existence is quite remarkable because, differently from all previous cases where $E_{3}$ occurred, both populations explode. A large herd of preys moves in our artificial space, hunted by an even larger herd of predators. Preys reproduce while running away from predators, predators reproduce while chasing preys generating the two moving columns illustrate in Figure 2.
@@
One may argue that humans are the only species able to make predictions based on variations of the populations of other species, and that since humans are predators, only the case 10 occurs in the real world. Specifically, humans can achieve co-existence with other species by observing their variations and behaving accordingly. This is interesting in itself, but we already mentioned that the Lotka-Volterra model has many interpretations besides the purely ecological one.
In particular, let us consider the application of the Lotka-Volterra model to the rather irregular oscillations of economic activities with a period of about 8-10 years, known as business cycle or trade cycle [35] [55] [78]. This tradition goes back to Marx, who remarked that capitalists cannot make profits if after many years of economic growth unemployment has become so low and salaries so high that capitalists do not make a profit by running their companies. This circumstance plunges the economy into recession, until unemployment becomes sufficiently high to induce workers to accept lower salaries again [60]. The Lotka-Volterra model can capture this dynamics by identifying capitalists with predators and workers with preys [35] [55] [78].
However, the aggregate Lotka-Volterra generates business cycles without economic growth. By contrast, our agent-based Lotka-Volterra combines growth with the business cycle. This aspect of our model deserves some comments.
Capitalism is — among else — a way of thinking, based on postponing consumption in order to save and invest. It requires making extrapolations about the future, a circumstance which, according to our computational experiments, turns oscillations around a fixed point into oscillations around a growing average. Notably, our model suggests that economic growth can only set in once the vast majority of actors have acquired an extrapolation-based way of thinking, a circumstance which in pre-modern societies may have never realized. Albeit kings did order the construction of infrastructures that would benefit society at large in the long run, insofar most people reasoned in terms of enjoying life day-by-day (carpe diem), indefinite economic growth could not start.
@@ 4 Conclusions
Since we investigated such ubiquitous dynamics as those expressed by preys and predators, one is tempted to ask whether our findings have some validity beyond this specific model. In particular, one may sensibly ask whether there exists a wider class of models where individual prediction by linear extrapolation generates interesting collective dynamics. And, one may also extend this question from the first to the second derivative and so forth, asking whether higher-order derivatives at the individual level are key to generate novel collective dynamics.
Tentatively, we are inclined to answer the first question with a “Yes” upon condition that the vast majority of individuals make predictions, whereas we bet that the answer to the second question is a clear “No.” Our rationale is that if most individuals compute first-order derivatives, then they can impact on one another generating higher-order derivatives. This is likely to happen in any model where interactions generate multiplicative, snowball-like effects that end up with constructing higher-order derivatives at the collective level. By contrast, higher-order derivatives at the individual level would add very little to the possibilities for interesting collective dynamics and, if anything, they could generate overfitting and individual complex dynamics cancelling out. These are pure speculations but, in a nutshell, the idea is that distributed systems are capable of doing sophisticated computations if individuals provide them with the basic bricks, namely, first-order derivatives.
The discovery of a coexistence-with-explosion dynamics that emerges in the prey-predator model if all agents make linear predictions highlights the presence of conflicting values in science, as well as in society at large. While everybody agrees that co-existence of species is a good thing, explosive dynamics is generally valued very differently by economists (GDP growth) and environmentalists (non-sustainable anthropic activities).
Evidently, the difference is in implicit predictions. Environmentalists implicitly predict that explosive dynamics will eventually bring about generalized extinction, whereas the economists do not. Both positions are tenable, precisely because they are implicit. Since this sort of long-run dynamics are not in the model, whether co-existence plus explosive dynamics represents an instance of collective intelligence is, within this model, an undecidable question.
@@
The authors declare not to have any conflict of interest whatsoever with the organizations, institutions and persons directly or indirectly mentioned in this paper. The authors did not receive any funding to carry out this research.
The authors received permission to reproduce the images that appear in Figure 1 from their copyright holders. All other figures and tables have been generated by the authors. No empirical data have been used.
Authors are listed in alphabetical order. GF contributed the theoretical architecture and literature review, AP contributed the computational model and exploration of results. The behavioural choices of artificial agents have been thoroughly discussed by both authors.
@@ Appendix A Lyapunov’s Stability Theorem and its Application to the Lotka-Volterra Model
Let us check whether $V(x,y)>0$. Let us consider the first term, namely, $d(x-x^{*}\ln x)$. It is $d(x-x^{*}\ln x)>0$ if $dx-c\ln x>0$rightarrow $dx>c\ln x$rightarrow $x>c/d\ln x$rightarrow $e^{x}>e^{c/d\ln x}$rightarrow $e^{x}>x^{c/d}$ which is always true because the exponential function yields greater values than the power function. The third term, $d(x^{*}-x^{*}\ln x^{*})$, makes sure that the combination of the first and the third term starts to yield positive values just as soon as $V$ leaves equilibrium $(x^{*},y^{*})$. Likewise, $e^{y}>y^{a/b}$ is always true because the exponential function yields greater values than the power function and the fourth term $b(y^{*}-y^{*}\ln y^{*})$ ensures that this only happens outside $(x^{*},y^{*})$. Thus, condition (2) is satisfied.
Let us turn to the third condition. It is $\dot{V}$ $=$ $(d-d\frac{x^{*}}{x})(ax-bxy)+(b-b\frac{y^{*}}{y})(-cy+dxy)$ $=$ $ad(x-x^{*})-bc(y-y^{*})+bd(x^{*}y-y^{*}x)$. Obviously, $\dot{V}(x^{*},y^{*})=0$. Let us investigate which parameters make for $\dot{V}(x,y)<0$ at $(x,y)\neq(x^{*},y^{*})$.
@@ Appendix B Our Agent-Based Lotka-Volterra
Stop-rep-if-wolves$>$sheep Off Off Off Off On Off On Off Off Off Off Off Off
Stop-rep-if-sheep$>$wolves Off Off Off Off Off On On Off Off Off Off Off Off
@@
In order to observe outputs we also introduced a parameter Transitory and a parameter Benchmarking that we set at 50 and 100 steps, respectively. Between Transitory and Benchmarking the model determines which dynamics count as oscillations. At each step within this interval, the minimum and maximum number of sheep and wolves that have been ever attained are updated. Let us call them $MINs$, $MAXS$, $MINW$ AND $MAXW$, respectively. Sheep oscillations are counted if $(MAXS-MINS)/2\geq 0.2(MAXS+MINS)/2$. If this condition is satisfied, one oscillation occurs if the number of sheep is greater or equal to $(MAXS+MINS)/2+0.5(MAXS-MINS)/2)$ and, subsequently, the number of wolves is smaller or equal to $(MAXW+MINW)/2)-0.5(MAXW-MINW)/2$. Wolves oscillations are counted in a similar way. Also the possibility of perfect foresight is only considered after Benchmarking steps have elapsed.

### v5

@@ 3.1 It’s Difficult to Leave a KISS
As mentioned above, the basic Wilensky and Reisman’s model does not support $E_{3}$. Suppose that predators and preys devise a reproductive strategy which, in their intentions, should be able to keep the system at $E_{3}$. With a slight departure from pure KISS, they may imagine that a flexible and adaptive strategy based on some negative feedback may generate co-existence. For instance, in order not to be either too many when the other species shrinks, or too few when the other species thrives, either predators, or preys, or both of them may decide to bind their reproduction to the other species’s success:
@@
Thus, this experiment apparently confirms Proposition 1, namely that individual intelligence destroys collective intelligence. However, these results were obtained with individuals who were not even attempting to figure out the global consequences of their behavior. It was a very slight departure from pure KISS. Perhaps, meta-cognition — the maximum KIDS of Proposition 3 — can help devising individual behavior that generates co-existence.
3.2 The Good KIDS
Let us explore the possibility that, as expressed by Proposition 3, at least some individuals can envision collective dynamics and therefore are able to influence it by selecting some appropriate behavior. Specifically, let us suppose that either predators, or preys, or both of them take on a collaborative attitude that may favor a collective agreement as prescribed by Qualification 3.3.
@@ 3.2 The Good KIDS
Finally, these behaviors can be combined by assuming that both predators and preys stop reproducing if their own population becomes larger than the other one. While the previous decisions were the outcome of agreements between either predators or preys, the combined behavior would be the outcome of a general agreement involving both populations.
@@
If predators behave as in 4, $E_{1}$ is reached only $2.8\%$ of the times, $E_{2}$ just $0.5\%$ of the times, whereas $E_{3}$ is reached $96.7\%$ of the times. Thus, this behavior succeeds to make predators and preys co-exist. To a closer scrutiny, the successful $96.7\%$ arises out of a $43.7\%$ of outcomes where the two populations are roughly constant, a $13.2\%$ where the population of preys oscillates whereas predators do not, a $21.0\%$ where the population of predators oscillates whereas preys do not and a $18.8\%$ where both populations oscillate.
@@
If both predators and preys take on the collaborative attitude 6 the outcome is somewhat mixed, with $E_{1}$ nonexistent at $0.0\%$ whereas both $E_{2}$ and $E_{3}$ are quite substantial at $22.9\%$ and $77.1\%$, respectively. Co-existence is still the most likely outcome but happens less often than in 4. Unlimited growth of preys is substantial but far from inevitable as in case 5.
Here it is worth noticing that, in the Lotka-Volterra model, predators and preys do not enjoy the same status. On the one hand, predators who are capable of meta-cognition are in position to steer the ecology towards co-existence by exerting double-loop learning (Proposition 3.1). On the other hand, preys capable of meta-cognition understand the system but cannot steer it, as it is typical of deutero-learning (Proposition 3.2). Preys can obtain dominance in the short run, but at the cost of unsustainable explosive dynamics for themselves.
@@
The problem with case 4 is that it requires predators to sacrifice their immediate interests in order to achieve co-existence. In practice, this may be difficult to attain.
3.3 The Perfect KIDS
One other direction for going into KIDS is that of combining Proposition 2 with Proposition 3 to obtain agents who have the ability to make predictions based on global understanding of the Lotka-Volterra model that they inhabit. Specifically, predators predict their own extinction if they either see the population of preys or their own population growing very quickly, whereas preys predict their own extinction if they see predators growing very quickly.
If they are rational, neither predators nor preys wait for their Lotka-Volterra world to unfold their inevitable destiny. As Qualification 2 suggests, they collectively agree to commit suicide if they see their inevitable end forthcoming.
Notably, readiness to commit suicide does not necessarily imply that this extreme decision is really carried out. Thus, let us explore the following behavioral rules:
@@ 3.3 The Perfect KIDS
Rules 7, 8 and 9 require an additional parameter to specify what “much faster” means. I introduced this parameter with the default assumption that “much faster” means ten times faster, supporting this assumption with a sensitivity analysis reported in Appendix C.
@@
In the end, by assuming that either predators or preys, or both of them are capable to predict the correct and inevitable outcome of the Lotka-Volterra model, co-existence $E_{3}$ can never be reached. Also in this case, KISS (Proposition 1) was better than KIDS.
@@ 3.4 Prediction based on Extrapolation
Similarly to§ 3.3, also in this case it is necessary to add a parameter in order to specify how many past steps the above variations are computed on. I set this parameter to the default value of five steps, supporting this choice with a sensitivity analysis reported in Appendix C.
Behaviour 10 is the second most effective in generating co-existence after 4 (altruistic predators seeking a collective agreement, § 3.2). In particular, with 10 generalized extinction $E_{1}$ never occurs. Preys-dominated equilibrium $E_{2}$ is reached $8.4\%$ of the times but, most importantly, $E_{3}$ is reached $91.6\%$ of the times. Out of this $91.6\%$, in $7.7\%$ both populations oscillate, $78.1\%$ only predators oscillate, in just $0.4\%$ of cases the opposite happens and, finally, in the remaining $5.4\%$ neither population oscillates. In a nutshell, the most likely outcome is that the population of predators oscillates while the population of preys remains roughly constant.
@@
Most interestingly, the mixed arrangement 12 is not some sort of average between 10 and 11 but rather a case on its own. With 12, $E_{1}$ is reached $17.9\%$ of the times, $E_{2}$ a mere $2.4\%$ whereas $E_{3}$ is reached $79.7\%$ of the times. Out of this $79,7\%$, $61.7\%$ of the times both populations oscillate, $1.4\%$ only the population of predators oscillates, $14.2\%$ the opposite happens and, finally, in the remaining $2.4\%$ neither population oscillates. Thus, co-existence is reached quite often, and most of the times it is reached with both populations oscillating.
This sort of co-existence is quite remarkable because, differently from all previous cases where $E_{3}$ occurred, both populations explode. A large herd of preys moves in artificial space, hunted by an even larger herd of predators. Preys reproduce while running away from predators, predators reproduce while chasing preys generating the two moving columns illustrate in Figure 2.
@@
One may argue that humans are the only species able to make predictions based on variations of the populations of other species, and that since humans are predators, only the case 10 occurs in the real world. Specifically, humans can achieve co-existence with other species by observing their variations and behaving accordingly. This is interesting in itself, but the Lotka-Volterra model has many interpretations besides the purely ecological one.
In particular, let us consider the application of the Lotka-Volterra model to the rather irregular oscillations of economic activities with a period of about 8-10 years, known as business cycle or trade cycle [34] [54] [78]. One explanation goes back to Marx, who remarked that after many years of economic growth unemployment has become so low and salaries so high that capitalists do not make a profit by running their companies. This circumstance plunges the economy into recession, until unemployment becomes sufficiently high to induce workers to accept lower salaries again [59]. The Lotka-Volterra model can capture this dynamics by identifying capitalists with predators and workers with preys [34] [54] [78].
However, the aggregate Lotka-Volterra generates business cycles without economic growth. By contrast, the agent-based Lotka-Volterra combines growth with the business cycle. This aspect deserves some comment.
Capitalism is — among else — a way of thinking, based on postponing consumption in order to save and invest. It requires making extrapolations about the future, a circumstance which, according to the above results, turns oscillations around a fixed point into oscillations around a growing average. Thus, the previous findings suggest that economic growth can only set in once the vast majority of actors have acquired an extrapolation-based way of thinking, a circumstance which in pre-modern societies may have never been realized. Albeit the kings of pre-capitalistic societies did order the construction of infrastructures that would benefit society at large in the long run, insofar most people reasoned in terms of carpe diem, indefinite economic growth could not start.
@@ 4 Conclusions
The exploration of individual algorithms more sophisticated than those expressed in the basic Lotka-Volterra model to generate collectively more “intelligent” behavior was in general unsuccessful, except for individual algorithms based on linear extrapolation. In a landscape that generally favors KISS over KIDS, the ability of single agents to make simple predictions stands out as a game-changer, at least insofar it concerns the Lotka-Volterra model.
Understandably, one is tempted to ask whether what similar individual behavior generates similar aggregate dynamics in more general models, or even in all models. In particular, one may sensibly ask whether there exists a wider class of models where individual prediction by linear extrapolation generates explosive and non-trivial collective dynamics. And, by extending this question from the first to the second derivative and so forth, one may ask whether higher-order derivatives at the individual level are key to generate novel collective dynamics.
Tentatively, I am inclined to answer the first question with a “Yes” upon condition that the vast majority of individuals make predictions, whereas I bet that the answer to the second question is a clear “No.” The rationale behind this stance is that if most individuals compute first-order derivatives, then they can impact on one another generating higher-order derivatives. This is likely to happen in any model where interactions generate multiplicative, snowball-like effects that end up with constructing higher-order derivatives at the collective level. By contrast, higher-order derivatives at the individual level would add very little to the possibilities for interesting collective dynamics and, if anything, they could generate overfitting or even cancel out individually complex dynamics. In a sense, this amounts to accepting very moderate KIDS, very close to the KISS principle, indeed.
The discovery of a coexistence-with-explosion dynamics that emerges in the prey-predator model if all agents make linear predictions highlights the presence of conflicting values in science, as well as in society at large. While everybody agrees that co-existence of species is a good thing, explosive dynamics is generally valued very differently by economists (GDP growth) and environmentalists (non-sustainable anthropic activities). Given so deeply-running differences, it not obvious whether we should label the explosive co-existence of species as a sort of “collective intelligence.”
Evidently, the difference is in implicit predictions. At the bottom line, environmentalists implicitly predict that explosive dynamics will eventually bring about generalized extinction, whereas economists implicitly predict that technological innovations will be able to curb undesired effects. Both positions are tenable because both are unproven, and both can not be proven also because they are implicit. Since this sort of long-run dynamics are not in the (however augmented) Lotka-Volterra model, whether explosive co-existence represents an instance of collective intelligence is, within this model, an undecidable question.
@@
The author declares not to have any conflict of interest whatsoever with the organizations, institutions and persons directly or indirectly mentioned in this paper. The author did not receive any funding to carry out this research.
The author received permission to reproduce the images that appear in Figure 1 from their copyright holders. All other figures and tables have been generated by the authors. No empirical data have been used.
@@ Appendix A Lyapunov’s Stability Theorem and its Application to the Lotka-Volterra Model
Let us check whether $V(x,y)>0$. Let us consider the first term, namely, $d(x-x^{*}\ln x)$. It is $d(x-x^{*}\ln x)>0$ if $dx-c\ln x>0$ $\rightarrow$ $dx>c\ln x$ $\rightarrow$ $x>c/d\ln x$ $\rightarrow$ $e^{x}>e^{c/d\ln x}$ $\rightarrow$ $e^{x}>x^{c/d}$ which is always true because the exponential function yields greater values than the power function. The third term, $d(x^{*}-x^{*}\ln x^{*})$, makes sure that the combination of the first and the third term starts to yield positive values just as soon as $V$ leaves equilibrium $(x^{*},y^{*})$. Likewise, $e^{y}>y^{a/b}$ is always true because the exponential function yields greater values than the power function and the fourth term $b(y^{*}-y^{*}\ln y^{*})$ ensures that this only happens outside $(x^{*},y^{*})$. Thus, condition (2) is satisfied.
Let us turn to the third condition. It is $\dot{V}$ == $(d-d\frac{x^{*}}{x})(ax-bxy)+(b-b\frac{y^{*}}{y})(-cy+dxy)$ == $ad(x-x^{*})-bc(y-y^{*})+bd(x^{*}y-y^{*}x)$. Obviously, $\dot{V}(x^{*},y^{*})=0$. Let us investigate which parameters make for $\dot{V}(x,y)<0$ at $(x,y)\neq(x^{*},y^{*})$.
@@ Appendix B Our Agent-Based Lotka-Volterra
Stop-rep-if-wolves>>sheep Off Off Off Off On Off On Off Off Off Off Off Off
Stop-rep-if-sheep>>wolves Off Off Off Off Off On On Off Off Off Off Off Off
@@
In order to observe outputs we also introduced a parameter Transitory and a parameter Benchmarking that we set at 50 and 100 steps, respectively. Between Transitory and Benchmarking the model determines which dynamics count as oscillations. At each step within this interval, the minimum and maximum number of sheep and wolves that have been ever attained are updated. Let us call them $MINs$, MAXSMAXS, MINWMINW AND $MAXW$, respectively. Sheep oscillations are counted if $(MAXS-MINS)/2\geq 0.2(MAXS+MINS)/2$. If this condition is satisfied, one oscillation occurs if the number of sheep is greater or equal to (MAXS+MINS)/2+0.5(MAXSMINS)/2)(MAXS+MINS)/2+0.5(MAXS-MINS)/2) and, subsequently, the number of wolves is smaller or equal to (MAXW+MINW)/2)0.5(MAXWMINW)/2(MAXW+MINW)/2)-0.5(MAXW-MINW)/2. Wolves oscillations are counted in a similar way. Also the possibility of perfect foresight is only considered after Benchmarking steps have elapsed.